3. We are given three coins. One has heads on both faces, the second has tails on both faces, and the third coin has a head on one face and a tail on the other face. We choose one coin at random, toss it, and observe that the result is heads. What is the probability that the opposite face is tails? 4, A survey shows that 40% of the resident of a particular town read the morning paper, 65% read the afternoon paper, and 10% read both. a. One resident is randomly selected. What is the probability that this person reads b. One resident is randomly selected. What is the probability that this person reads neither the morning nor the afternoon paper? only the morning paper?

Please show your work on 3&4

Gotta understand them really well.
Answer #1

3)P(heads appears)=P(head on both side and head appears)+P(head on one side and head appears)+P(tail on both side and head appears)=(1/3)*(1)+(1/3)*(1/2)+(1/3)*0=1/2

therefore P(tail on other side given heads appears)

=P(head on one side and head appears)/P(heads appears)=(1/3)*(1/2)/(1/2) =1/3

4)

a)P(read neither)=1-P(read at least one)

=1-(P(read morning paper)+P(read afternoon paper)-P(read both))

=1-(0.4+0.65-0.1)=0.05

b)

P(only morning paper)=P(read morning paper)-P(read both)=0.4-0.1=0.30